# Internal structure of the $T_{cc}(3875)^{+}$ from its light-quark mass dependence

### v1

@@ 4 Theoretical uncertainty and cross checks
Figure 7: Theoretical uncertainty for the $T_{cc}$ pole position. The bands represent the truncation uncertainty of the chiral expansion for the $T_{cc}$ pole position at $\Lambda=700$ MeV when the ${\cal O}(Q^{4})$ terms in Eq. (14) are included. The green band corresponds to the variation of the dimensionless constant $\alpha_{4}$ in the range $[-1,1]$, while the blue band corresponds to the variation of the dimensionless constant $\tilde{\alpha}_{4}$ in the same range. Red solid and dashed lines represent the best fits for $\Lambda=700$ MeV and $\Lambda=500$ MeV, respectively.
@@
$V_{\text{cont}}^{(4)}=D_{4}(\xi^{2}-1)(p^{2}+{p^{\prime}}^{2})+\tilde{D}_{4}(\xi^{4}-1)$ (14)
while the effect of $m_{\pi}$-independent $\mathcal{O}(p^{4})$ contact term is expected to be small and not considered. Here, $D_{4}$ and $\tilde{D}_{4}$ can be expressed as
@@
with $\Lambda_{\chi}\simeq 1$ GeV being the chiral symmetry breaking scale. Here, $\alpha_{4}$ and $\tilde{\alpha}_{4}$ are dimensionless prefactors expected to be of the order of $1$ based on naturalness, as discussed analogously in the $NN$ case [59]. Indeed, applying the same logic to the terms $C_{2}$ and $D_{2}$ in Eq. (2), we can rewrite them in the following form
@@
To estimate the impact of the $\mathcal{O}(Q^{4})$ terms, we allow $\alpha_{4}$ and $\tilde{\alpha}_{4}$ to vary from $-1$ to $+1$. This range is considered conservative, given that the values of $\{\alpha_{0},\alpha_{2},\tilde{\alpha}_{2}\}$ are smaller. The uncertainty associated with the truncation of the chiral expansion is shown in Fig. 7 (see the green and blue bands). To obtain these bands, we use the best fit parameters for $C_{0},C_{2}$ and $D_{2}$ from our NLO results and supplement the calculations with the higher-order potential $V_{\text{cont}}^{(4)}$ from Eq. (14). No refit of the phase shifts at $m_{\pi}=280$ MeV was performed, resulting in a more conservative estimate. It is reassuring that the spread in the results at $m_{\pi}=280$ MeV is quite natural, and it appears to be comparable in size to the statistical uncertainty from Fig. 5. Also, the change in the pole position at the physical pion mass is minor and lies well within the uncertainty estimated in Ref. [13]. The resulting uncertainty grows with $\xi$, but it remains comparable with the statistical error in Fig. 5.
@@ 4.1 Comparison to other works
Here, we briefly comment on other lattice calculations available in the literature and compare them with our results. In Ref. [21], elastic $DD^{*}$ scattering in S-wave was investigated. However, the study was limited by using only a single lattice volume and having only one data point in the near-threshold region. As a result, no pole position was extracted. In Ref. [24], the HAL QCD method was utilized to extract the $DD^{*}$ scattering potential at the pion mass $m_{\pi}=146.4$ MeV, which was then employed to calculate the phase shifts above the $DD^{*}$ threshold. No visible signature of the OPE was observed. Employing the ERE, the $T_{cc}$ pole was reported to be a virtual state with $k=(-8\pm 8^{+3}_{-5})i$ MeV corresponding to $E_{\rm pole}=-59\begin{subarray}{c}{}^{+53\,+2}\\ {}_{-99\,-67}\end{subarray}$ keV, where the errors represent statistical and systematic uncertainties, in order. Extrapolation of the HAL QCD potential to the physical pion mass, neglecting three-body and isospin violating effects, led to a bound state but with a binding energy substantially smaller than observed experimentally.
@@
Very recently, a lattice investigation of coupled $DD^{*}-D^{*}D^{*}$ scattering at $m_{\pi}=391$ MeV was presented in Ref. [22], finding a sizable coupled-channel effect. Using the Lüscher method and parameterizations of the coupled-channel amplitudes that ignore the OPE, two states were found in the energy plane: a virtual state about $62(34)$ MeV below the $DD^{*}$ threshold and a resonance about ($249(35)+i11(13)/$ 2) MeV below the $D^{*}D^{*}$ threshold. By employing a contact theory, we can indeed obtain a virtual state around $80$ MeV below the $DD^{*}$ threshold. However, this state transforms into a resonance with a substantially different pole position when the lhc effect from the OPE is included. Further investigations are needed to investigate the impact of the lhc on the coupled-channel transitions and the pole below the $D^{*}D^{*}$ threshold.
@@ 5 Summary and conclusions
The resulting position of the pole transitions from a quasi-bound to a bound, a virtual and a resonance state as the pion mass increases, indicating the molecular nature of the $T_{cc}$. To observe the effect of the OPE on the pole for pion masses slightly above the physical value may require very precise lattice simulations. However, for $m_{\pi}\gtrsim 230$ MeV, the presence of the OPE is evident as the repulsion it generates shifts the pole into the complex plane, leading to substantially different results from those obtained without pions.

### v2

@@ 4 Theoretical uncertainty and cross checks
4.1 Chiral truncation error
Figure 7: Theoretical uncertainty for the $T_{cc}$ pole position. The bands represent the truncation uncertainty of the chiral expansion for the $T_{cc}$ pole position at $\Lambda=700$ MeV when the ${\cal O}(Q^{4})$ terms in Eq. (14) are included. The green band corresponds to the variation of the dimensionless constant $\alpha_{4}$ in the range $[-1,1]$, while the blue band corresponds to the variation of the dimensionless constant $\tilde{\alpha}_{4}$ in the same range. Red solid and black dashed lines represent the best fits for $\Lambda=700$ MeV and $\Lambda=500$ MeV, respectively.
@@ 4.1 Chiral truncation error
$V_{\text{cont}}^{(4)}=D_{4}(\xi^{2}-1)(p^{2}+{p^{\prime}}^{2})+\tilde{D}_{4}(\xi^{4}-1),$ (14)
while the effect of the $m_{\pi}$-independent $\mathcal{O}(p^{4})$ contact term is neglected, as this is consistent with available experimental information at $m_{\pi}=m_{\pi}^{\text{ph}}$ and lattice data at $m_{\pi}=280$ MeV. Here, $D_{4}$ and $\tilde{D}_{4}$ can be expressed as
@@
with $\Lambda_{\chi}\simeq 1$ GeV being the chiral symmetry breaking scale. Here, $\alpha_{4}$ and $\tilde{\alpha}_{4}$ are dimensionless prefactors expected to be of the order of $1$ based on naturalness, as discussed analogously in the $NN$ case [66]. Indeed, applying the same logic to the terms $C_{2}$ and $D_{2}$ in Eq. (2), we can rewrite them in the following form,
@@
To estimate the impact of the $\mathcal{O}(Q^{4})$ terms, we allow $\alpha_{4}$ and $\tilde{\alpha}_{4}$ to vary from $-1$ to $+1$. This range is considered conservative, given that the values of $\{\alpha_{0},\alpha_{2},\tilde{\alpha}_{2}\}$ are smaller. The uncertainty associated with the truncation of the chiral expansion is shown in Fig. 7 (see the green and blue bands). To obtain these bands, we use the best fit parameters for $C_{0},C_{2}$ and $D_{2}$ from our NLO results and supplement the calculations with the higher-order potential $V_{\text{cont}}^{(4)}$ from Eq. (14). No refit of the phase shifts at $m_{\pi}=280$ MeV was performed, resulting in a more conservative estimate. It is reassuring that the spread in the results at $m_{\pi}=280$ MeV is quite natural, and it appears to be comparable in size to the statistical uncertainty from Fig. 5. The resulting uncertainty grows with $\xi$, but it remains comparable with the statistical error in Fig. 5.
@@
Additionally, we verified that the change in $\text{Im}E_{\text{pole}}$ of the $T_{cc}$ at the physical pion mass, due to the inclusion of the contact term $C_{2}$, with this LEC fixed at $m_{\pi}=280$ MeV, is very small after adjusting $C_{0}$ to reproduce $\text{Re}E_{\text{pole}}$, and remains well within the uncertainty estimated in Ref. [17].
Finally, another potential source of uncertainty arises from the transition between the real world and the isospin limit at the physical pion mass. As discussed above, the parameter $C_{0}$ was adjusted to reproduce the real part of the $T_{cc}$ pole position, $\text{Re}E_{\text{pole}}=-356$ keV, as extracted from the coupled-channel analysis in Ref. [17]. However, in the isospin limit, the $T_{cc}$ becomes slightly more bound. To estimate the impact of this effect on the pole trajectory, we adopt the following strategy: (i) Starting from the contact coupled-channel framework of Ref. [17] (see fit 1 in Tables I and II), we use the extracted value of $C_{0}$ to calculate the $T_{cc}$ pole in the isospin limit, which gives $\text{Re}E_{\text{pole}}\approx-940$ keV; (ii) Relying on the fact that $\text{Re}E_{\text{pole}}$ at the physical pion mass is largely insensitive to pion dynamics [17], we use this binding energy as input in the full pionful framework in the isospin limit, recalculate $C_{0}$ [^2] Note that we cannot directly use $C_{0}$ from the coupled-channel pionful fit of [17] in the isospin limit for the following reasons: (i) unlike Du et al., we employ a formalism with relativistic pions; (ii) in the full approach with pions, $C_{0}$ is known to exhibit a limit cycle, meaning it can rapidly change from $\infty$- ∞ to +$\infty$+ ∞ as a function of the cutoff [67, 68] while the observables remain cutoff independent. Due to these reasons, taking $C_{0}$ from one pionful framework and applying it in a different one could lead to large, uncontrolled systematic errors. This, however, is expected to work well in the contact case., and find that the impact of this effect on the pole trajectory is very minor (just slightly larger than the effect of cutoff variation shown in Fig. 7). As expected for a more attractive potential, all transitions (quasi-bound to bound, to virtual, to resonance) occur at slightly larger pion masses.
@@ 4.2 Comparison to other works
Here, we briefly comment on other lattice calculations available in the literature and compare them with our results. In Ref. [28], the HAL QCD method was utilized to extract the $DD^{*}$ scattering potential at the pion mass $m_{\pi}=146.4$ MeV, which was then employed to calculate the phase shifts above the $DD^{*}$ threshold. No visible signature of the OPE was observed. Employing the ERE, the $T_{cc}$ pole was reported to be a virtual state with $k=(-8\pm 8^{+3}_{-5})i$ MeV corresponding to $E_{\rm pole}=-59\begin{subarray}{c}{}^{+53\,+2}\\ {}_{-99\,-67}\end{subarray}$ keV, where the errors represent statistical and systematic uncertainties, in order. Extrapolation of the HAL QCD potential to the physical pion mass, neglecting three-body and isospin violating effects, led to a bound state but with a binding energy substantially smaller than observed experimentally.
@@
In Ref. [25], elastic $DD^{*}$ scattering in S-wave was investigated at $m_{\pi}\approx 348.5$ MeV. However, the study was limited by using only a single lattice volume and having only one data point in the near-threshold region. As a result of the ERE fits to the phase shifts obtained using the Lüscher method, the effective range parameters were presented – the scattering length and the effective range. Using these parameters, we extracted the pole which turns out to be a virtual state about $-28^{+4}_{-5}$ MeV, as shown in Fig. 6 by the green dot. While this result ignores the physics related with long-range interactions, it appears consistent with our prediction based on contact EFT.
Very recently, a lattice investigation of coupled $DD^{*}-D^{*}D^{*}$ scattering at $m_{\pi}=391$ MeV was presented in Ref. [26], finding a sizable coupled-channel effect. Using the Lüscher method and parameterizations of the coupled-channel amplitudes that ignore the OPE, two states were found in the energy plane: a virtual state about $62(34)$ MeV below the $DD^{*}$ threshold and a resonance about ($249(35)+i11(13)/$ 2) MeV below the $D^{*}D^{*}$ threshold. By employing a contact theory, we can indeed obtain a virtual state around $75$ MeV below the $DD^{*}$ threshold. However, this state transforms into a resonance with a substantially different pole position when the lhc effect from the OPE is included. Further investigations are needed to investigate the impact of the lhc on the coupled-channel transitions and the pole below the $D^{*}D^{*}$ threshold.
@@ 5 Summary and conclusions
The resulting position of the pole transitions from a quasi-bound to a bound, a virtual and a resonance state as the pion mass increases, indicating the molecular nature of the $T_{cc}$. Observing the effect of the OPE on the pole for pion masses slightly above the physical value is challenging and may require very precise lattice simulations. However, for $m_{\pi}\gtrsim 230$ MeV, the presence of the OPE is evident as the repulsion it generates shifts the pole into the complex plane, leading to substantially different results from those obtained without pions.